OpenStudy (anonymous):

For matinees,a movie theater charges $7 per adult and $4 per child.if 84 people attended a matinee and paid a total of $432,How many adults and how many children were there .

OpenStudy (anonymous):

where x= adults and y = children eq 1: 432 = 7*x + 4*y and eq 2: 84 = x+y set y = 84-x, then substitute this y into equation 1

OpenStudy (anonymous):

can you solve this for x? 432 = 7*x + 4*(84-x)

OpenStudy (anonymous):

lets assume all were child then charge=84*4=$336 but he got $432 432-336=$96 extra and 7-4=$3 per head extra 96/3=32 adults remaining child